jowhite:

Change the unit of length. 1yd 1 ft = _____ yd

jowhite:

snowflake0531:

1 yard + 1 foot = ___ yards 3 feet = 1 yards 1 foot = \(\frac{1}{3}\) yards 1 yard + 1 foot = 1 yard + \(\frac{1}{3}\) yards = 1\(\frac{1}{3}\) yards